find the slope of any line parallel to the line through points (15,1) and (4,2)
Answer by venugopalramana(3286)

find the slope of any line parallel to the line through points A....X1,Y1(15,1) and B..X2,Y2(4,2)
SLOPE OF PARALLEL LINE = SLOPE OF LINE JOINING THESE 2 POINTS A & B.=(Y2-Y1)/(X2-X1)=(2-1)/(4-15)=-1/11